# Teaching Transformers to Solve Combinatorial Problems through Efficient Trial & Error

### v1

@@ 4 Beyond imitation learning – optimizing guesses
Our method is illustrated in Figure 5. Unlike our earlier deterministic (argmax) generation policy, here we sample guesses from the model’s token distribution at the guess node. If the sampled guess leads to a full solution via rule application, the process terminates; otherwise, it restarts. The guess node, viewed as a single decision step, may represent either a single action/token or a thinking sequence (see Appendix D). This process continues until a valid trajectory terminates at an “end” token, indicating completion. This framework offers a tractable yet powerful setting to study how well LLMs can prioritize high-value guesses under uncertainty.
@@
Central to our optimization is the training objective applied at the output of the guess node, which must encourage the model to concentrate probability mass on all valid backdoor moves. Specifically, if $S$ is the set of valid backdoor moves at a given guess node, and $p_{i}$ denotes the probability assigned to each such move, we consider different losses for optimization. The loss corresponding to the expected number of trials is $\mathcal{L}_{1}=(\sum_{i\in S}p_{i})^{-1}$ whereas the standard Cross-Entropy loss with multiple targets is $\mathcal{L}_{2}=-\sum_{i\in S}\log p_{i}$. For numerical stability during training, we minimize $\log\mathcal{L}_{1}$, implemented using the log-sum-exp trick. One can interpret the L1 loss in terms of a geometric distribution: we treat each good guess as a Bernoulli trial where a “success” means finding a valid backdoor. Not every correct guess actually works as a backdoor, so each attempt can succeed or fail. The goal is to minimize the expected number of trials needed to find a working backdoor; analogous to minimizing the expectation in the geometric distribution.

### v2

@@ 4 Beyond imitation learning – optimizing guesses
Our method is illustrated in Figure 5. Unlike our earlier deterministic (argmax) generation policy, here we sample guesses from the model’s token distribution at the guess node. If the sampled guess leads to a full solution via rule application, the process terminates; otherwise, it restarts. Crucially, only successful trajectories, those containing the correct guess, are used for training. This design makes the approach more efficient than our previous imitation learning setup: while imitation learning includes incorrect guess tokens and thus provides richer supervision, it also encourages the model to learn to predict subsequent wrong tokens via next-token prediction, rather than immediately identifying the correct path. By excluding incorrect guesses from the training data, this approach goes beyond imitation learning focuses learning on high-value decisions. The guess node, viewed as a single decision step, may represent either a single action/token or a thinking sequence (see Appendix D). This process continues until a valid trajectory terminates at an “end” token, indicating completion. This framework offers a tractable yet powerful setting to study how well LLMs can prioritize high-value guesses under uncertainty.
@@
Central to our optimization is the training objective applied at the output of the guess node, which must encourage the model to concentrate probability mass on all valid backdoor moves. Specifically, if $S$ is the set of valid backdoor moves at a given guess node, and $p_{i}$ denotes the probability assigned to each such move, we consider different losses for optimization. The loss corresponding to the expected number of trials is $\mathcal{L}_{1}=(\sum_{i\in S}p_{i})^{-1}$ whereas the standard Cross-Entropy loss with multiple targets is $\mathcal{L}_{2}=-\sum_{i\in S}\log p_{i}$. For numerical stability during training, we minimize $\log\mathcal{L}_{1}$, implemented using the log-sum-exp trick. One can interpret the $\mathcal{L}_{1}$ loss in terms of a geometric distribution: we treat each good guess as a Bernoulli trial where a “success” means finding a valid backdoor. Not every correct guess actually works as a backdoor, so each attempt can succeed or fail. The goal is to minimize the expected number of trials needed to find a working backdoor; analogous to minimizing the expectation in the geometric distribution.